Hulme Company operates a small manufacturing facility as a supplement to its regular service activities. At the beginning of 2020, an asset account for the company showed the following balances:

Manufacturing equipment $ 132,000
Accumulated depreciation through 2019 76,000

During 2020, the following expenditures were incurred for the equipment:

Major overhaul of the equipment on January 2, 2020, that improved efficiency $ 18,000
Routine maintenance and repairs on the equipment 2,600


The equipment is being depreciated on a straight-line basis over an estimated life of 15 years with a $18,000 estimated residual value. The annual accounting period ends on December 31.

Prepare the adjusting entry that should be made by Hulme Company at the end of 2020 for depreciation of the manufacturing equipment, assuming no change in the original estimated life or residual value.
Answer #1
Date General Journal Debit Credit
December 31,2020 Depreciation expense 11200
     Accumulated Depreciation-Equipment 11200
Workings:
Cost of Equipment 132000
Less: Residual value 18000
Depreciable cost 114000
Divide by useful life 15
Annual Depreciation 7600
Years of Equipment used 10 =76000/7600
Cost of Equipment 150000 =132000+18000
Less: Accumulated depreciation through 2019 76000
Book value at the beginning of 2020 74000
Less: Residual value 18000
Remaining Depreciable cost 56000
Divide by remaining useful life 5 =15-10
Annual Depreciation 11200